EA011A

PI: Alef
Title: Cyg A

Session: Oct 1996 (session 3)
Frequency: 6 cm
Mode: A
Obs time (UT): 300 EU 10:30(26/10)-22:30(26/10)
Stations: EbJbNtOnToShSmUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (J. Conway, A.Polatidis at Mon Oct 28 10:22:10 EST 1996) Success
No known problems. High winds

Torun (Roman Feiler at Mon Oct 28 19:09:42 EST 1996) Success
No known problems. Only left polarisation recorded.

Westerbork (tony foley at Tue Oct 29 09:07:56 EST 1996) Success
Looks OK, and no extra attenuation was needed

Noto (Corrado Trigilio at Wed Oct 30 05:55:32 EST 1996) Minor failures
One scan lost for cable wrap problems (scan 3 of tape 3). No other problems reported. Covered sky.

Shanghai (Huang Xinyong at Tue Nov 5 03:49:40 EST 1996) Success
No problems. Clock offset 0.0 micro sec. Weather was cloudy.

Jodrell (Alastair Gunn at Thu Nov 7 07:04:37 EST 1996) Success
At 20:51 UT telescope was not on source since source change had to be done manually. VC attenuation changed. Tapes shipped to Bonn on 28/10/96 via Woodcroft.

Urumqi (Liu Xiang at Mon Nov 18 08:56:21 EST 1996) Minor failures
Clock offset 0.2+/-0.4us.

Simeiz (Alexandre Volvatch at Tue Nov 19 03:54:04 EST 1996) Minor failures
6 passes (20:51 - 22:07) were omited because we had not power (switch off).

MPIfR (Walter Alef at Tue Sep 23 17:14:55 METDST 1997) Minor failures
Westerbork had only linear polarization and was therefore not correlated. 5 of the remaining 8 stations had only 1 polarization. Rest O.K.